One thousand tickets were sold at a game. Adult tickets cost $8.50 while kids cost $4.50, for a total of $7,300. How many tickets of each kind were sold?
Answer by mananth(16946)

One thousand tickets were sold at a game. Adult tickets cost $8.50 while kids cost $4.50, for a total of $7,300. How many tickets of each kind were sold?
..
adult tickets ----------- x numbers -----------8.50
kids tickets--------------y numbers------------4.50
total 7300
..
x+y =1000.....................1
8.5x+4.5y= 7300...............2
...
multiply (1) by -8.5
-8.5x-8.5y = -8500
..
add this to (2)
-4x=-1200
/4
x= 300 adult tickets
plug value of x in(1)
y =700 kids tickets.
.
m.ananth@hotmail.ca